OpenStudy (anonymous):

How long will a plane flying at a speed of 145 miles per hour take to travel a distance of 870 miles?

OpenStudy (neer2890):

Distance= speed* time so time= distance/ speed time= 870/145 =6 hours

OpenStudy (neer2890):

and 145=870/x i think @aminamoazzam you have made a mistake in this x=145*870 ??

OpenStudy (aminamoazzam):

yeah got it !!!